# One size doesn’t fit all: Predicting the Number of Examples for In-Context Learning

### v2

@@ Abstract
In-context learning (ICL) refers to the process of adding a small number of localized examples (ones that are semantically similar to the input) from a training set of labelled data to an LLM’s prompt with an objective to effectively control the generative process seeking to improve the downstream task performance. Existing ICL approaches use an identical number of examples (a pre-configured hyper-parameter) for each data instance. Our work alleviates the limitations of this ‘one fits all’ approach by dynamically predicting the number of examples for each data instance to be used in few-shot inference with LLMs. In particular, we employ a multi-label classifier, the parameters of which are fitted using a training set, where the label for each instance in the training set indicates if using a specific value of $k$ (number of most similar examples from 0 up to a maximum value) leads to correct $k$-shot downstream predictions. Our experiments on a number of text classification benchmarks show that AICL substantially outperforms standard ICL by up to 17%.
One size doesn’t fit all: Predicting the Number of Examples for In-Context Learning
Manish Chandra, Debasis Ganguly  and Iadh Ounis University of Glasgow Glasgow, United Kingdom m.chandra.1@research.gla.ac.uk, Debasis.Ganguly@glasgow.ac.uk, iadh.ounis@glasgow.ac.uk
@@ 1 Introduction
Large Language Models (LLMs) exhibit remarkable abilities to model text semantics in an abstract and general manner without any task specific training Radford et al. (2018). In-context learning (ICL)111also interchangeably known as few-shot learning or retrieval-augmented generation (RAG) with ground-truth labels makes use of this abstract representation and knowledge representation capabilities of LLMs Brown et al. (2020); Arora et al. (2022); Weidinger et al. (2022) to address a range of different downstream tasks. For instance, such models can provide effective solutions for tasks including assessing reviews Mysore et al. (2023), answering questions Li et al. (2023a), recommending relevant documents Pradeep et al. (2023) etc., by using only a small number of (even zero) examples without any task-specific training.
@@
Figure 1: a) Example workflow of ICL for sentiment classification. The illustrative example shows a test instance for which a single demonstration (as retrieved from the training set) does not result in a correct prediction (prediction workflow of the red arrows). The example also shows that increasing the number of demonstrations from one to two results in a correct prediction (green arrows). We propose a method to estimate the number of examples that are likely to yield a correct prediction. b) Motivation behind using a variable number of examples for ICL across the test instances: 1) The test sample ‘?’ is located within a homogeneous neighborhood of negative data points indicating that an LLM may perform well with only a few nearest neighbour demonstrations. 2) The test instance ‘?’ is located within a heterogeneous neighborhood, as a result of which, an LLM may require a higher number of such demonstrations to correctly predict its class.
@@
More formally, ICL refers to the process of conditioning an LLM’s decoder (frozen parameters) towards generating potentially relevant outputs that could then be interpreted in the context of a specific task, e.g., words such as ‘great’ and ‘wow’ can be mapped to a positive sentiment Schick et al. (2020). The output of an LLM’s decoder is controlled by varying the input to the LLM, which is usually structured in the form of an instruction (the task description), and a small number of representative examples Ni et al. (2022). Figure 1(a) shows an example of how ICL works for the downstream task of movie review sentiment analysis. It can be seen that both demonstrations (shown in blue) and the current instance form a part of the input to an LLM. Researchers have worked towards adapting an ICL workflow in several ways, which range from using localised examples Liu et al. (2022a) instead of random ones, to diversifying these examples Levy et al. (2023), and also ideally ordering them Kumar and Talukdar (2021); Rubin et al. (2022). What prior work has not addressed so far is the effect of dynamically selecting the number of examples in an ICL workflow. Our idea takes motivation from information retrieval (IR), where different queries exhibit different levels of retrieval performance mainly due to the inherent characteristic of the information need itself Kanoulas et al. (2011), or how well-formulated the query is Datta et al. (2022). Drawing the parallel that a query in IR is analogous to a test instance in ICL and that the localised examples are potentially relevant documents Rubin et al. (2022), we hypothesise that some test instances are associated with better candidates for training examples (i.e., examples which are useful in the sense that including them as a part of the prompt leads to correct predictions), and hence including a small number of these examples should be adequate. On the other hand, the retrieval quality for some test instances used as queries do not yield good candidates, as a result of which, one needs to further look down the ranked list to collect the useful ones Bahri et al. (2020); Ganguly and Yilmaz (2023). The parallel with IR means that the notion of relevance of a document to a query needs to be replaced by the downstream usefulness of an example to a test instance.
@@
The idea of choosing a variable number of localised examples is also somewhat similar to selecting a variable-sized neighborhood for $k$-NN classification Zhong et al. (2017), where the key idea is that a homogeneous neighborhood is likely to require a relatively small-sized neighborhood for a correct prediction, and a heterogeneous one would likely require a larger one (see Figure 1(b)). The same idea can be applied to ICL, where a test instance that is similar to a number of training instances with conflicting labels may require a larger number of examples.
@@
In our work, we apply a supervised learning based workflow to learn a mapping between the features of a test instance (embedding plus the label distribution of its neighborhood) and the ideal number of examples that should be used in ICL for the downstream prediction. More specifically, for each instance of the training set we vary $k$ the number of ICL examples - within a range of 00 to a pre-configured maximum value (say $M$) and store each indicator of whether a particular value of $k$ leads to a correct prediction as an $(M+1)$-length Boolean vector (e.g., the indicator for $k=1$ in Figure 1(a) is negative, whereas the one for $k=2$ is positive). We then train a multi-label classifier on these pairs of instances and the Boolean indicators, the assumption being that similar instances with similar label distributions should also exhibit a similar distribution over the values of $k$ leading to correct $k$-shot learning. During inference time, for each test instance, we apply the multi-label classifier to obtain a candidate list of predicted $k$ values, and we select the one for which the prediction confidence is the highest. We then use these many examples for the ICL-based downstream prediction. This means that as per the schematics of Figure 1(a), we can potentially find out a ‘green path’ leading to a correct prediction for each test instance among several other ‘red paths’ that lead to incorrect ones.
@@ 2 Related Work
LLMs, when scaled from millions to billions of parameters, have been demonstrated to be adaptable to a broad set of tasks due to instruction tuning Ouyang et al. (2022); Brown et al. (2020), in the sense that they are not only able to produce semantically correct and coherent text, but are also able to adapt themselves surprisingly well with small changes in contexts supplied as inputs, commonly called prompts Arora et al. (2022). Previous research studies the problem of constructing an appropriate prompt for LLMs from two broad perspectives: a) prompt tuning in the embedding space Li and Liang (2021); Liu et al. (2022b); Qin and Eisner (2021); Liu et al. (2023), and b) prompt searching in the text space Lu et al. (2022); Zhang et al. (2023); Diao et al. (2023); Liu et al. (2022a); Shi et al. (2023). Prompt tuning is a lightweight alternative to fine-tuning, which keeps language model parameters frozen and instead optimizes a sequence of continuous task-specific vectors. The key idea of prompt tuning is to inject task-specific embedding into hidden layers and then tune these embeddings using gradient-based optimization. However, these methods require the modification of the original inference process of the model, which is impractical for the case of black-box LLM services, such as GPT3 Brown et al. (2020). Furthermore, prompt tuning introduces additional computational and storage costs, which is typically expensive for LLMs. A more efficient way is to optimize prompting via searching appropriate demonstration samples and ordering them in the original text space.
@@
ICL has advanced the use of LLMs for task-specific applications with minimal examples, forming a solid foundation for subsequent investigations into optimizing this learning paradigm Dong et al. (2022). In ICL, a small number of labeled examples from a training set are appended to a prompt instruction to control the text generation of LLM so that it is beneficial to a downstream task Mysore et al. (2023); Li et al. (2022); Ni et al. (2021); Pradeep et al. (2023). In addition to leveraging ICL for a purely generative task, e.g., question answering or abstractive summarisation Brown et al. (2020); Li et al. (2023a); Tang et al. (2023), a more common use is in a predictive task, such as text classification Lu et al. (2022); Milios et al. (2023); Wei et al. (2021), where each class is specified by a set of words, commonly called a verbaliser Schick and Schütze (2021). Once each class for a predictive task is well-defined, the generated text can be mapped to the most likely class by using the probabilities of the tokens generated by the decoder.
A crucial aspect of ICL is the adept selection and utilization of the demonstrations for task comprehension. Recent works have explored the dynamics of in-context examples, elucidating how ICL enhances the language model’s performance across various NLP tasks by providing a minimal set of examples at inference time Han et al. (2023). Furthermore, the effectiveness of ICL is significantly shaped by the strategies for selecting and ordering demonstrations. For instance, the work in Luo et al. (2023) expanded the applicability of retrieval-based ICL approaches by demonstrating that even simple word-overlap similarity measures such as BM25 outperform randomly selected demonstrations. Methods such as KATE Liu et al. (2022a) also showed that localised examples work better than random ones. Different from KATE that uses a fixed-size neighbourhood for each data instance, we propose to modify a standard ICL workflow by employing a variable number of examples.
@@ 3.1 Standard In-Context Learning (ICL)
In-context learning (ICL), unlike supervised learning, does not involve training a set of parameters on labeled examples. Rather, the posteriors are now a function of the following: a) text of the input test instance, b) the decoder parameters of a pre-trained LLM, c) a prompt instruction, and d) optionally, a set of $k$ input examples, commonly called $k$-shot prompting Liu et al. (2022a). Formally,
@@
where, different from a supervised setup, the function $f$ does not have a parameterized representation that can be learned using a training set with gradient descent. The function itself depends on the pre-trained frozen parameters $\phi_{\text{LLM}}$ of an LLM, the current inputs for which a label is to be predicted, and a prompt comprising a set of $k$ text units denoted by $\pazocal{N}_{k}(\mathbf{x})$. This set $\pazocal{N}_{k}(\mathbf{x})=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{k}\}$ of Equation 1 (where each $\mathbf{z}_{j}\in\pazocal{T}$, a training set of labelled data instances) in ICL is usually comprised of localised examples, i.e., examples that are topically similar to the current instance Liu et al. (2022a); Luo et al. (2024). Particularly for our experiments, we employ SBERT Reimers and Gurevych (2019) as the neighborhood similarity computation function as per the findings of Liu et al. (2022a).
@@ 3.2 Adaptive ICL
We now describe our methodology that uses a variable number of examples by extending the standard ICL workflow of Equation 1. We call our method ‘Adaptive In-Context Learning’, or AICL for short. The idea of AICL centres around choosing the context $\pazocal{N}_{k}(\mathbf{x})$ in a data-driven manner, i.e., making $k$ a function of the data, i.e., the current instance $\mathbf{x}$ itself. This is analogous to choosing a different value of $k$ for a $k$-NN based non-parametric model or choosing a different rank cut-off for top-retrieved set of documents for different queries Bahri et al. (2020); Ganguly and Yilmaz (2023). The motivation is that classifying some instances would be more difficult than others, in which cases they are potentially to benefit from a larger value of $k$ (more context). On the other hand, for relatively easier data instances, using too much context may be detrimental for an effective prediction.
@@
For each training set instance $\mathbf{x}\in\pazocal{T}$, we first execute $k$-shot inference with progressively increasing values of $k$ within $\{0,\ldots,M\}$, where $M$ is a pre-configured threshold (specifically, $M=10$ in our experiments). This is depicted by the progressively increasing colored dotted regions within the ‘classifier training’ block of Figure 2. After obtaining $M+1$ posteriors $P(\hat{y}|\mathbf{x})$ for an instance $\mathbf{x}$ of the training set, we construct an $(M+1)$ dimensional Boolean vector, the $i^{th}$ component of which indicates if using $i$ ICL examples lead to a correct downstream prediction for the current instance $\mathbf{x}$ (note that since $\mathbf{x}$ is a training set instance, the knowledge of its true label $y(\mathbf{x})$ is available). Formally speaking,
@@
The multi-label classifier $\theta:\mathbf{x}\mapsto\pazocal{K}(\mathbf{x})$ means that the number of ICL examples depend only on the document content, or in other words, topically similar content potentially requires a similar number of ICL examples for accurate predictions. However, this does not take into account the class label distribution (the $y(\mathbf{x})$ values) of the examples. In fact, it is reasonable to assume that an instance with a more homogeneous neighborhood (in terms of the class label distribution) likely indicates an easier instance and hence may require a smaller number of examples, whereas a more heterogeneous neighborhood may indicate a larger number of examples (e.g., see Figure 1(b)).
This hypothesis means that the use of neighborhood class distribution as additional features to encode each training data instance ($\mathbf{x}\in\pazocal{T}$) is likely to lead to a better predictor $\theta$, which in turn, is likely to lead to better estimations of $\kappa(\mathbf{x})$ for AICL. Specifically, to incorporate the class prior information to aid predicting the number of ICL examples, we append an $M$ dimensional vector $\mathbf{l}\in\mathbb{Z}_{p}^{M}$ ($p$ is the number of class labels) to each SBERT embedded input vector $\mathbf{x}\in\pazocal{T}$. The $i^{th}$ component of this vector contains the class label of the $i^{th}$ neighbor in $\pazocal{N}_{M}(\mathbf{x})$, i.e., $\mathbf{l}=\{y(\mathbf{z}_{1}),\ldots,y(\mathbf{z}_{M})\}$. With this modification, the multi-label classifier now learns:
$\theta:\mathbf{x}\oplus\mathbf{l}\mapsto\pazocal{K}(\mathbf{x}),$ (6)
where $\mathbf{l}$ denotes the class label distribution of $M$ nearest neighbors, and $\oplus$ denotes the concatenation operator.

### v3

@@ Abstract
In-context learning (ICL) refers to the process of adding a small number of localized examples from a training set of labelled data to an LLM’s prompt with an objective to effectively control the generative process seeking to improve the downstream task performance. Existing ICL approaches use an identical number of examples (a pre-configured hyper-parameter) for each data instance. Our work alleviates the limitations of this ‘one fits all’ approach by dynamically predicting the number of examples for each data instance to be used in few-shot inference with LLMs. In particular, we employ a multi-label classifier, the parameters of which are fitted using a training set, where the label for each instance in this training set indicates if using a specific value of $k$ (number of most similar examples from 0 up to a maximum value) leads to correct $k$-shot downstream predictions. Our experiments on a number of text classification benchmarks show that AICL substantially outperforms standard ICL by up to 17%.
Keywords:
Large Language Model Adaptive In-context Learning Multi-label Classification
@@ 1 Introduction
Large Language Models (LLMs) exhibit remarkable abilities to model text semantics in an abstract and general manner without any task specific training [35]. In-context learning (ICL) [^1] Also interchangeably known as few-shot learning or retrieval-augmented generation (RAG) with ground-truth labels. makes use of this abstract representation and knowledge representation capabilities of LLMs [3, 1, 45] to address a range of different downstream tasks. For instance, such models can provide effective solutions for tasks including assessing reviews [28], answering questions [15], recommending relevant documents [33] etc., by using only a small number of (even zero) examples without any task-specific training.
@@
Figure 1: a) Example workflow of ICL for sentiment classification: The example shows a test instance for which a single demonstration (as retrieved from the training set) does not result in a correct prediction (prediction workflow of the red arrows). It also shows that increasing the number of demonstrations from one to two results in a correct prediction (green arrows). We propose a method to estimate the number of examples that are likely to yield a correct prediction. b) Motivation behind using a variable number of examples for ICL across the test instances: The test sample ‘?’ is located within a homogeneous neighborhood of negative data points indicating that an LLM may perform well with only a few nearest neighbour demonstrations. The test instance ‘?’, on the other hand, is located within a heterogeneous neighborhood, as a result of which, an LLM may require a higher number of such demonstrations to correctly predict its class.
@@
More formally, ICL refers to the process of conditioning an LLM’s decoder (frozen parameters) towards generating potentially relevant outputs that could then be interpreted in the context of a specific task, e.g., words such as ‘great’ and ‘wow’ can be mapped to a positive sentiment [38]. The output of an LLM’s decoder is controlled by varying the input to the LLM, which is usually structured in the form of an instruction (the task description), and a small number of representative examples [29]. Figure 1(a) shows an example of how ICL works for the downstream task of movie review sentiment analysis. It can be seen that both demonstrations (shown in blue) and the current instance form a part of the input to an LLM. Researchers have worked towards adapting an ICL workflow in several ways, which range from using localised examples [19] instead of random ones, to diversifying these examples [13], and also ideally ordering them [12, 37]. What prior work has not addressed so far is the effect of dynamically selecting the number of examples in an ICL workflow. Our idea takes motivation from information retrieval (IR), where different queries exhibit different levels of retrieval performance mainly due to the inherent characteristic of the information need itself [11], or how well-formulated the query is [5]. Drawing the parallel that a query in IR is analogous to a test instance in ICL [32] and that the localised examples are potentially relevant documents [37], we hypothesise that some test instances are associated with better candidates for training examples (i.e., examples which are useful in the sense that including them as a part of the prompt leads to correct predictions), and hence including a small number of these examples should be adequate. On the other hand, the retrieval quality for some test instances used as queries do not yield good candidates, as a result of which, one needs to further look down the ranked list to collect the useful ones [2, 8]. The parallel with IR means that the notion of relevance of a document to a query needs to be replaced by the downstream usefulness of an example to a test instance.
@@
The idea of choosing a variable number of localised examples is also somewhat similar to selecting a variable-sized neighborhood for $k$-NN classification [48], where the key idea is that a homogeneous neighborhood is likely to require a relatively small-sized neighborhood for a correct prediction, and a heterogeneous one would likely require a larger one (see Figure 1(b)). The same idea can be applied to ICL, where a test instance that is similar to a number of training instances with conflicting labels may require a larger number of examples.
@@
In our work, we apply a supervised learning based workflow to learn a mapping between the features of a test instance (embedding plus the label distribution of its neighborhood) and the ideal number of examples that should be used in ICL for the downstream prediction. More specifically, for each instance of the training set, we vary $k$ the number of ICL examples - within a range of 00 to a pre-configured maximum value (say $M$) and store each indicator of whether a particular value of $k$ leads to a correct prediction as an $(M+1)$-length Boolean vector (e.g., the indicator for $k=1$ in Figure 1(a) is negative, whereas the one for $k=2$ is positive). We then train a multi-label classifier on these pairs of instances and the Boolean indicators, the assumption being that similar instances with similar label distributions should also exhibit a similar distribution over the values of $k$ leading to correct $k$-shot learning. During inference time, for each test instance, we apply the multi-label classifier to obtain a candidate list of predicted $k$ values, and we select the one for which the prediction confidence is the highest. We then use these many examples for the ICL-based downstream prediction. This means that as per the schematics of Figure 1(a), we can potentially find out a ‘green path’ leading to a correct prediction for each test instance among several other ‘red paths’ that lead to incorrect ones.
@@ 2 Related Work
LLMs, when scaled from millions to billions of parameters, have been demonstrated to be adaptable to a broad set of tasks due to instruction tuning [31, 3], in the sense that they are not only able to produce semantically correct and coherent text, but are also able to adapt themselves surprisingly well with small changes in contexts supplied as inputs, commonly called prompts [1]. Previous research studies the problem of constructing an appropriate prompt for LLMs from two broad perspectives: a) prompt tuning in the embedding space [16, 20, 34, 21], and b) prompt searching in the text space [22, 47, 6, 19, 40]. Prompt tuning is a lightweight alternative to fine-tuning, which keeps language model parameters frozen and instead optimizes a sequence of continuous task-specific vectors. The key idea of prompt tuning is to inject task-specific embedding into hidden layers and then tune these embeddings using gradient-based optimization. However, these methods require the modification of the original inference process of the model, which is impractical for the case of black-box LLM services, such as GPT3 [3]. Furthermore, prompt tuning introduces additional computational and storage costs, which is typically expensive for LLMs. A more efficient way is to optimize prompting via searching appropriate demonstration samples and ordering them in the original text space.
@@
ICL has advanced the use of LLMs for task-specific applications with minimal examples, forming a solid foundation for subsequent investigations into optimizing this learning paradigm [7]. In ICL, a small number of labeled examples from a training set are appended to a prompt instruction to control the text generation of LLM so that it is beneficial to a downstream task [28, 14, 30, 33]. In addition to leveraging ICL for a purely generative task, e.g., question answering or abstractive summarisation [3, 15, 41], a more common use is in a predictive task, such as text classification [22, 26, 44], where each class is specified by a set of words, commonly called a verbaliser [39]. Once each class for a predictive task is well-defined, the generated text can be mapped to the most likely class by using the probabilities of the tokens generated by the decoder.
A crucial aspect of ICL is the adept selection and utilization of the demonstrations for task comprehension. Recent works have explored the dynamics of in-context examples, elucidating how ICL enhances the language model’s performance across various NLP tasks by providing a minimal set of examples at inference time [10]. Furthermore, the effectiveness of ICL is significantly shaped by the strategies for selecting and ordering demonstrations. For instance, the work in [23] expanded the applicability of retrieval-based ICL approaches by demonstrating that even simple word-overlap similarity measures such as BM25 outperform randomly selected demonstrations. Methods such as KATE [19] also showed that localised examples work better than random ones. Different from KATE that uses a fixed-size neighbourhood for each data instance, we propose to modify the standard ICL workflow by employing a variable number of examples.
@@ 3.1 Standard In-Context Learning (ICL)
In-context learning (ICL), unlike supervised learning, does not involve training a set of parameters on labeled examples. Rather, the posteriors are now a function of the following: a) text of the input test instance, b) the decoder parameters of a pre-trained LLM, c) a prompt instruction, and d) optionally, a set of $k$ input examples, commonly called $k$-shot prompting [19]. Formally,
@@
where, different from a supervised setup, the function $f$ does not have a parameterized representation that can be learned using a training set with gradient descent. The function itself depends on the pre-trained frozen parameters $\phi_{\text{LLM}}$ of an LLM, the current inputs for which a label is to be predicted, and a prompt comprising a set of $k$ text units denoted by $\pazocal{N}_{k}(\mathbf{x})$. This set $\pazocal{N}_{k}(\mathbf{x})=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{k}\}$ of Equation 1 (where each $\mathbf{z}_{j}\in\pazocal{T}$, a training set of labelled data instances) in ICL is usually comprised of localised examples, i.e., examples that are topically similar to the current instance [19, 24]. Particularly for our experiments, we employ SBERT [36] as the neighborhood similarity computation function as per the findings of [19].
@@ 3.2 Adaptive ICL
We now describe our methodology that uses a variable number of examples by extending the standard ICL workflow of Equation 1. We call our method ‘Adaptive In-Context Learning’, or AICL in short. The idea of AICL centres around choosing the context $\pazocal{N}_{k}(\mathbf{x})$ in a data-driven manner, i.e., making $k$ a function of the data, i.e., the current instance $\mathbf{x}$ itself. This is analogous to choosing a different value of $k$ for a $k$-NN based non-parametric model or choosing a different rank cut-off for top-retrieved set of documents for different queries [2, 8]. The motivation is that classifying some instances would be more difficult than others, in which cases they are potentially to benefit from a larger value of $k$ (more context). On the other hand, for relatively easier data instances, using too much context may be detrimental for an effective prediction.
@@
For each training set instance $\mathbf{x}\in\pazocal{T}$, we first execute $k$-shot inference with progressively increasing values of $k$ within $\{0,\ldots,M\}$, where $M$ is a pre-configured threshold (specifically, $M=10$ in our experiments). This is depicted by the progressively increasing colored dotted regions within the ‘classifier training’ block of Figure 2. After obtaining $M+1$ posteriors $P(\hat{y}|\mathbf{x})$ for an instance $\mathbf{x}$ of the training set, we construct an $(M+1)$ dimensional Boolean vector, the $i^{th}$ component of which indicates if using an $i$ number of ICL examples lead to a correct downstream prediction for the current instance $\mathbf{x}$ (note that since $\mathbf{x}$ is a training set instance, the knowledge of its true label $y(\mathbf{x})$ is available). Formally speaking,
@@
The multi-label classifier $\theta:\mathbf{x}\mapsto\pazocal{K}(\mathbf{x})$ means that the number of ICL examples depend only on the document content, or in other words, topically similar content potentially requires a similar number of ICL examples for accurate predictions. However, this does not take into account the class label distribution, i.e., the $y(\mathbf{x})$ values of the examples. In fact, it is reasonable to assume that an instance with a more homogeneous neighborhood (in terms of the class label distribution) likely indicates an easier instance and hence may require a smaller number of examples, whereas a more heterogeneous neighborhood may indicate a larger number of examples (e.g., see Figure 1(b)).
This hypothesis means that the use of neighborhood class distribution as additional features to encode each training data instance ($\mathbf{x}\in\pazocal{T}$) is likely to lead to a better predictor $\theta$, which in turn, is likely to lead to better estimations of $\kappa(\mathbf{x})$ for AICL. Specifically, to incorporate the class prior information to aid predicting the number of ICL examples, we append an $M$ dimensional vector $\vec{l}\in\mathbb{Z}_{p}^{M}$ ($p$ is the number of class labels) to each SBERT embedded input vector $\mathbf{x}\in\pazocal{T}$. The $i^{th}$ component of this vector contains the class label of the $i^{th}$ neighbor in $\pazocal{N}_{M}(\mathbf{x})$, i.e., $\vec{l}=\{y(\mathbf{z}_{1}),\ldots,y(\mathbf{z}_{M})\}$. With this modification, the multi-label classifier now learns:
$\theta:\mathbf{x}\oplus\vec{l}\mapsto\pazocal{K}(\mathbf{x}),$ (6)
where $\vec{l}$ denotes the class label distribution of $M$ nearest neighbors, and $\oplus$ denotes the concatenation operator.